# Non-degenerate mixed maps and contact structures

### v1

@@ 2.1. Mixed maps
(2.1) $\begin{split}\alpha_{1}(t)\left(\left(\frac{\partial f^{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\right)^{I}_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{1}-p_{i}}+o(t)\right)+&\dots+\alpha_{k}(t)\left(\left(\frac{\partial f^{k}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}\right)_{P}^{I}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{k}-p_{i}}+o(t)\right)\\ &=\\ {}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}(t)\left(\left({}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{1}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}}\right)^{I}_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{1}-p_{i}}+o(t)\right)+&\dots+{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}(t)\left(\left({}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{k}}{\partial z_{i}}}\right)^{I}_{P}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{k}-p_{i}}+o(t)\right),\end{split}$
We may suppose $q_{1}+d_{1}\leq\dots\leq q_{k}+d_{k}$. Let $J$ denote the indices $j=1,\dots,l$, with $l\leq k$, for which $q_{1}+d_{1}=\dots=q_{l}+d_{l}$. Comparing the orders of both sides, it follows that the differentials $\left({}\mkern 3.0mu\overline{\mkern-3.0muDf^{1}}\right)^{I}_{P},\dots,\left({}\mkern 3.0mu\overline{\mkern-3.0muDf^{k}}\right)^{I}_{P}$, $\left({}\mkern 3.0mu\overline{\mkern-3.0muD}f^{1}\right)^{I}_{P},\dots,\left({}\mkern 3.0mu\overline{\mkern-3.0muD}f^{k}\right)^{I}_{P}$ are linearly dependent at $a\in\left(F^{I}_{P}\right)^{-1}(0)\cap\mathbb{C}^{*m}$, where we set zero for coefficients whose indices do not belong to $J$. This leads to a contradiction with the partial non-degeneracy of $F$. ∎
@@ 2.2. Mixed Hamm Complete Intersections
A particular example of genuine mixed map defining an ICIS can be obtained from the previous constructions as follows. Let $\textbf{a}=(a_{1},\dots,a_{n}),$ be a vector of positive integers and $\Lambda=(\lambda_{ij})$ a complex matrix of order $n\times k$. For each $i=1,\dots,k$, let $f^{i}=\sum_{j=1}^{n}\lambda_{ij}z_{j}^{a_{j}}$ be a complex Pham-Brieskorn polynomial. Hamm showed in [14] that for a sufficiently general matrix, the map germ $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ defines an ICIS, which we shall refer to as Hamm ICIS.
Let $\textbf{b}=(b_{1},\dots,b_{n})$ be a second vector of non negative integers and consider the mixed Pham-Brieskorn polynomial $g^{i}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{j=1}^{n}\lambda_{ij}z_{i}^{a_{i}+b_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}^{b_{i}}$. This type of mixed functions is a particular case of the construction discussed in Subsection 1.2. If we require all $k\times k$-minors of $\Lambda$ being nonzero, the map $H:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ whose coordinate functions are $h^{i}(z)=\sum_{j=1}^{n}\lambda_{ij}z_{j}$ is non-degenerate. By Proposition 2.9 and Theorem 2.10 we conclude that the mixed Hamm map defines an ICIS, which we call mixed Hamm ICIS. Notice that $G$ is not an algebraic ICIS by Proposition 1.13.
In [33, Theorem 4.1] it is shown that complex and mixed Pham-Brieskorn polynomials are topologically equivalent, and this assertion easily applies for the maps constructed above. Furthermore, Oka proved in [27] that the links are smoothly equivalent. In this section, we extend this result for mixed Hamm ICIS. We fix the vectors a and b of integers. For each $i=1,\dots,k$, let $f^{i}$ and $g^{i}$ be mixed and complex Pham-Brieskorn polynomials, respectively, as before. Define the following family of mixed maps:
$G_{t}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=(1-t)G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})+tF(z,{}\mkern 3.0mu\overline{\mkern-3.0muz}),$
where $F=(f^{1},\dots,f^{k})$ and $G=(g^{1},\dots,g^{k})$. Let us denote $V_{t}^{i}=\left(g_{t}^{i}\right)^{-1}(0)$, where $g_{t}^{i}$ are the coordinate functions of $G_{t}$, and $V_{t}=G_{t}^{-1}(0)$. We fix the notation $\mathbb{C}^{n}_{*}:=\mathbb{C}^{n}\setminus\{0\}$.
@@
    The variety $V_{t}$ intersects the sphere $\mathbb{S}_{r}^{2n-1}$ transversely for any $r>0$.
@@
    $\psi_{t}:(\mathbb{B}_{r}^{2n},E_{0}(r))\longrightarrow(\mathbb{B}_{r}^{2n},E_{t}(r)),$
    where $E_{t}(r)=\{z\in\mathbb{C}^{n}_{*}:G_{t}(z)=0,\|z\|\leq r\}$. Moreover, it restricts also as diffeomorphisms
    $\psi_{t}:(\mathbb{S}_{r}^{2n-1},\partial E_{0}(r))\longrightarrow(\mathbb{S}_{r}^{2n-1},\partial E_{t}(r)),$
@@
In the first item, for $t=0$ and $t=1$ the assertion is already proved. In the other cases, it is enough to notice that the Newton polyhedron of $g_{t}^{i}$ is the polyhedron of a complex Pham-Brieskorn polynomial and we reduce to a case already settled. The second item is a consequence of [27, Lemma 2], which shows the assertion for each $V_{t}^{i}$ and thus for the intersection $V_{t}=V_{t}^{1}\cap\dots\cap V_{t}^{k}$. The last statement follows from Ehresmann fibration theorem for subbundles applied to the canonical projections
@@
Let $F$ and $G$ be the complex and mixed Hamm ICIS, respectively, and let $r>0$ be fixed.
@@
    $\psi:(\mathbb{S}_{r}^{2n-1},K_{G})\longrightarrow(\mathbb{S}_{r}^{2n-1},K_{F}),$
    where $K_{G}$ and $K_{F}$ are the links defined by $F$ and $G$, respectively.
@@
    Let $I\subset\{1,\dots,n\}$ such that $|I|>k$. Then the map $\psi$ also restricts to a diffeomorphism
    $\psi^{I}:(\mathbb{S}_{r}^{2|I|-1},K_{G^{I}})\longrightarrow(\mathbb{S}_{r}^{2|I|-1},K_{F^{I}}),$
    where $K_{G^{I}}$ and $K_{F^{I}}$ are the links defined by the restrictions $F^{I}$ and $G^{I}$, respectively.
2.3. Milnor fibrations

### v3

@@ 2.1. Mixed maps
(2.2) $\begin{split}\alpha_{1}(t)\left(\frac{\partial f^{1,I}_{P}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{1}-p_{i}}+o(t)\right)+&\dots+\alpha_{k}(t)\left(\frac{\partial f^{k,I}_{P}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{k}-p_{i}}+o(t)\right)\\ &=\\ -{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}(t)\left({}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{1,I}_{P}}{\partial z_{i}}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{1}-p_{i}}-o(t)\right)-&\dots-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}(t)\left({}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{k,I}_{P}}{\partial z_{i}}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{k}-p_{i}}+o(t)\right),\end{split}$
Comparing the lower degree terms of these equations for every $i=1,\dots,n$, the point $a\in(F^{I})_{P}^{-1}(0)\cap\mathbb{C}^{*I}$ becomes a critical point of $(F^{I})_{P}$. Given that $|I|>k$, one has $f^{j,I}\not\equiv 0$ for every $j$, and thus we may apply the first item of Proposition 2.9 to conclude that $F^{I}$ is also non-degenerate, which leads to a contradiction. ∎
Corollary 2.14.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a non-degenerate and convenient mixed map germ. Then $F^{I}:(\mathbb{C}^{I},0)\longrightarrow(\mathbb{C}^{k},0)$ is a mixed ICIS for every $I\subset\{1,\dots,n\}$ such that $|I|>k$ and $V_{F^{I}}$ has positive dimension.
Proof.
As we have seen, the convenience allows us to apply the first item of Proposition 2.9 to conclude that $F^{I}$ is non-degenerate. For the convenience itself, let $J\subset I$. First, if $|J|=k$, then:
$V_{F^{I}}\cap\mathbb{C}^{J}=V_{F}\cap\mathbb{C}^{I}\cap\mathbb{C}^{J}=V_{F}\cap\mathbb{C}^{J}=\{0\}.$
If $|J|>k$, then $\left(f^{j,I}\right)^{J}=f^{j,I\cap J}\not\equiv 0$, since $|I\cap J|=|J|>k$. Therefore, the result follows from Theorem 2.13. ∎
Definition 2.15.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a convenient mixed map germ. We say that $F$ is strongly convenient if, for every vector $P$ of positive integers, the face map $F_{P}:(\mathbb{C}^{I},0)\longrightarrow(\mathbb{C}^{k},0)$, where $I\subset\{1,\dots,n\}$ is a non-empty subset, satisfies the following conditions:
  1. (1)
    If $|I|>k$, then $F_{P}$ is convenient.
  2. (2)
    If $|I|\leq k$, then $V_{F_{P}}=0$.
If we look at the complete convenience of all face functions for $k=1$, the strong convenience means that every compact face intersects all coordinate axes of the hyperplanes containing it.
Example 2.16.
Let $f(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{i=1}^{n}z_{i}^{a_{i}+b_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}^{b_{i}}$ be a mixed Pham-Brieskorn polynomial, where $a_{i},b_{i}\geq 0$ are integers. The face functions of $f(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ are the restrictions of $f$ to the subspaces $\mathbb{C}^{I}$, and thus it is strongly convenient. More generally, let $f(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ be a mixed function whose Newton polyhedron coincides with that of a Pham-Brieskorn polynomial, with terms possibly above the polyhedron or exactly in the compact faces determined by some subset of vertices in the coordinate axis. Then $f(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ is strongly convenient. For instance, $f(z)=z_{1}^{2}+z_{2}^{2}+z_{2}z_{3}+z_{3}^{2}+T(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$, whose face functions are $f$ itself, $z_{1}^{2},z_{2}^{2}$, $z_{1}^{2}+z_{2}^{2}$, or $z_{2}^{2}+z_{2}z_{3}+z_{3}^{2}$, provided that $T(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$ consists of terms above the Newton polyhedron.
Example 2.17.
Let $F=(f^{1},f^{2}):(\mathbb{C}^{4},0)\longrightarrow(\mathbb{C}^{2},0)$ be the mixed map germ whose Newton polyhedron coincides with that of the mixed map given by the coordinate functions $g^{1}=z_{1}^{a}+z_{2}^{b}+z_{3}^{2}+z_{3}^{2}z_{4}$ and $g^{2}=-z_{1}^{a}+2z_{2}^{b}-z_{3}^{2}+z_{4}^{2}$. The face functions do not involve the monomial $z_{2}^{3}z_{4}$, which lies on a non-compact face. Thus, we reduce the analysis to maps whose coordinate functions are Pham-Brieskorn polynomials. It is straightforward to check that the face maps $F_{P}$ are non-degenerate, and that these are convenient in the sense of Definition 2.12.
Moreover, in the next subsection, Example 2.25 presents another strongly convenient map germ. More generally, we can also consider the pullback of strongly convenient mixed maps by mixed coverings to produce more examples as follows.
Proposition 2.18.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a (strongly) convenient mixed map germ and $\phi$ a mixed covering. Then $G=\phi^{*}F$ is also (strongly) convenient.
Proof.
We first observe two facts. For each $I$, the coordinate functions are related as follows:
$g^{j,I}=\left(f^{j}\circ\phi\right)^{I}=f^{j}\circ\phi^{I}=f^{j,I}\circ\phi,$
since the coordinate functions of $\phi$ are in separated variables. This implies that $g^{j,I}\not\equiv 0$ if and only if $f^{j,I}\not\equiv 0$. Secondly, $\phi(V_{G})=V_{F}$. Suppose that there exists $0\neq p\in V_{G}\cap\mathbb{C}^{I}$, where $|I|=k$. Then $p\in\mathbb{C}^{*J}$, for some $J\subset I$. Since $\phi$ preserves $\mathbb{C}^{*J}$, we conclude that $\phi(p)\in V_{F}\cap\mathbb{C}^{*J}$. By Proposition 2.10, this implies that $\phi(p)=0$, and then $p=0$, which is a contradiction. For the strong convenience, we have seen that the face maps $G_{\tilde{P}}=F_{P}\circ\phi$, for vectors $P,\tilde{P}$ of positive integers. Since the assertion holds for each $F_{P}$, we conclude by the first part. ∎
Proposition 2.19.
Let $F:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a non-degenerate strongly convenient mixed map germ and $I\subset\{1,\dots,n\}$ such that $|I|>k$. Then, for every vector $P$ of positive integers such that $V_{F^{I}_{P}}$ has positive dimension, $F^{I}_{P}$ is a mixed ICIS.
Proof.
First, we have that $F^{I}$ is non-degenerate, since $F$ is convenient. On the other hand, since $f^{j,I}\not\equiv 0$ for every $j$, by the same argument of the first item of Proposition 2.9, $(F^{I})_{P}=F_{Q}$ for some vector $Q$. By the strong convenience, $F_{Q}$ is convenient. The result follows from Theorem 2.13. ∎
Lastly, according to [30, Lemma 28], strongly non-degenerate mixed functions have an isolated critical value. In the higher dimensional case, we also prove a similar result about critical points of $F$.
Proposition 2.20.
Let $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ be a strongly non-degenerate mixed map germ and $p\in\Sigma_{F}$. Then $f^{j}(p)=0$ for some $j=1,\dots,k$.
Proof.
We apply an argument similar to that in the proof of [30, Lemma 28]. As before, by the Curve Selection Lemma and Proposition 1.1, if the statement is false, there exist real analytic curves $w(t)\subset\Sigma_{F}$ and $\alpha_{1}(t),\dots,\alpha_{k}(t)\subset\mathbb{C}$, non-simultaneously vanishing, such that $f^{j}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))\neq 0$ for every $j$ and
$\alpha_{1}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{1}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)+\dots+\alpha_{k}(t){}\mkern 3.0mu\overline{\mkern-3.0muD}f^{k}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)=-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{1}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right)-\dots-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}(t){}\mkern 3.0mu\overline{\mkern-3.0muDf^{k}}\left(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t)\right).$
Let $I=\{i:w_{i}(t)\not\equiv 0\}$ and write the respective analytic expansions:
$f^{j,I}(w(t),{}\mkern 3.0mu\overline{\mkern-3.0muw}(t))$ $=c_{j}t^{r_{j}}+o(t),\;c_{j}\in\mathbb{C}^{*}\,,r_{j}\geq 1,$
$w_{i}(t)$ $=a_{i}t^{p_{i}}+o(t),\;a_{i}\in\mathbb{C}^{*}\,,p_{i}\geq 1,$
$\alpha_{j}(t)$ $=b_{j}t^{q_{j}}+o(t),\;b_{j}\in\mathbb{C}^{*}\,,q_{j}\geq 1,$
for all $i\in I$ and $j$ such that $\alpha_{j}(t)\not\equiv 0$. Let $P=(p_{1},\dots,p_{n})$ and $a=(a_{1},\dots,a_{n})$. For each $i$, the relation above becomes:
(2.3) $\begin{split}\alpha_{1}(t)\left(\frac{\partial f^{1,I}_{P}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{1}-p_{i}}+o(t)\right)+&\dots+\alpha_{k}(t)\left(\frac{\partial f^{k,I}_{P}}{\partial{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{k}-p_{i}}+o(t)\right)\\ &=\\ -{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{1}(t)\left({}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{1,I}_{P}}{\partial z_{i}}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{1}-p_{i}}-o(t)\right)-&\dots-{}\mkern 3.0mu\overline{\mkern-3.0mu\alpha}_{k}(t)\left({}\mkern 3.0mu\overline{\mkern-3.0mu\frac{\partial f^{k,I}_{P}}{\partial z_{i}}}(a,{}\mkern 3.0mu\overline{\mkern-3.0mua})t^{d_{k}-p_{i}}+o(t)\right),\end{split}$
where $d_{j}$ is the weighted homogeneous degree of the face function $f^{j,I}_{P}$. Equation 2.3 now implies that $a\in\mathbb{C}^{*I}$ is a critical point of $\left(F^{I}\right)_{P}$. As we proceeded in Proposition 2.9, since $f^{j,I}\not\equiv 0$ for every $j$, there exists a vector $Q$ such that $F^{I}_{P}=F_{Q}$. Thence, $(a,a^{\prime})\in\mathbb{C}^{*I}\times\mathbb{C}^{*I^{C}}$ is a critical point of $F_{Q}$ for every $a^{\prime}$, which contradicts the strong non-degeneracy of $F$. ∎
@@ 2.2. Mixed Hamm complete intersections
A particular example of a genuine mixed map defining an ICIS can be obtained from the previous constructions as follows. Let $\textbf{a}=(a_{1},\dots,a_{n}),$ be a vector of positive integers and $\Lambda=(\lambda_{ij})$ a complex matrix of order $n\times k$. For each $i=1,\dots,k$, let $f^{i}=\sum_{j=1}^{n}\lambda_{ij}z_{j}^{a_{j}}$ be a complex Pham-Brieskorn polynomial. Hamm showed in [16] that for a sufficiently general matrix, the map germ $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ defines an ICIS, which we shall refer to as Hamm ICIS.
Let $\textbf{b}=(b_{1},\dots,b_{n})$ be a second vector of non negative integers and consider the mixed Pham-Brieskorn polynomial $g^{i}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{j=1}^{n}\lambda_{ij}z_{i}^{a_{i}+b_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}^{b_{i}}$. This type of mixed function is a particular case of the construction discussed in Subsection 1.2. If we require all $k\times k$-minors of $\Lambda$ being nonzero, the map $H:(\mathbb{C}^{n},0)\longrightarrow(\mathbb{C}^{k},0)$ whose coordinate functions are $h^{i}(z)=\sum_{j=1}^{n}\lambda_{ij}z_{j}$ is non-degenerate. By Proposition 2.9 and Theorem 2.13, we conclude that the mixed Hamm map defines an ICIS, which we call mixed Hamm ICIS. If $b_{i}\geq 2$ for every $i=1,\dots,n$, notice that $G$ is not an algebraic ICIS by Proposition 1.13.
In [36, Theorem 4.1], it is shown that complex and mixed Pham-Brieskorn polynomials are topologically equivalent, and this assertion easily applies to the maps constructed above. Furthermore, Oka proved in [31] that the links are smoothly equivalent. In this section, we extend this result to mixed Hamm ICIS. We fix the vectors a and b of integers. For each $i=1,\dots,k$, let $f^{i}$ and $g^{i}$ be mixed and complex Pham-Brieskorn polynomials, respectively, as before. Define the following family of mixed maps:
(2.4) $G_{t}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=(1-t)G(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})+tF(z,{}\mkern 3.0mu\overline{\mkern-3.0muz}),$
where $F=(f^{1},\dots,f^{k})$ and $G=(g^{1},\dots,g^{k})$. Let us denote $V_{t}^{i}=\left(g_{t}^{i}\right)^{-1}(0)$, where $g_{t}^{i}$ are the coordinate functions of $G_{t}$, and $V_{t}=G_{t}^{-1}(0)$. We fix the notation $\mathbb{C}^{n}_{*}:=\mathbb{C}^{n}\setminus\{0\}$ and $\mathbb{B}_{*,r}^{2n}=\mathbb{B}_{r}^{2n}\setminus\{0\}$.
Remark 2.21.
Let $w\in\mathbb{C}^{n}_{*}$, $I=\{i:w_{i}=0\}$, and fix $t\in(0,1)$. In [31, Lemma 2], Oka proves the existence of a curve $\xi(r)=(\xi_{1}(r),\dots,\xi_{n}(r))\subset\mathbb{C}^{n}$ with the following properties:
  1. (1)
    The parameter $r\in(0,\infty)$;
  2. (2)
    For each $i\in I$, $\xi_{i}=0$;
  3. (3)
    For each $i$, $\xi_{i}^{a_{i}}(t+(1-t)\|\xi_{i}\|^{2b_{i}})=rw_{i}^{a_{i}}(t+(1-t)\|w_{i}\|^{2b_{i}})$;
  4. (4)
    The tangent vector $u=\frac{d\xi}{dr}(1)$ is not tangent to the sphere $\mathbb{S}_{r}^{2n-1}$.
Let $j=1,\dots,k$ and consider the coordinate function $g_{t}^{j}$. Suppose further that $w\in V_{t}^{i}\cap\mathbb{S}_{r}^{2n-1}$. Then $g_{t}^{j}(\xi,{}\mkern 3.0mu\overline{\mkern-3.0mu\xi})\equiv 0$ and $u$ is a tangent vector of $V_{t}^{i}$ at $w$ which is not tangent to the sphere. Hence, $V_{t}^{i}\pitchfork\mathbb{S}_{r}^{2n-1}$ for all $r>0$.
@@
    The variety $V_{t}$ intersects the sphere $\mathbb{S}_{r}^{2n-1}$ transversely for all $r>0$.
@@
    $\psi_{t}:(\mathbb{B}_{*,r}^{2n},E_{t}(r))\longrightarrow(\mathbb{B}_{*,r}^{2n},E_{1}(r)),$
    where $E_{t}(r)=\{z\in\mathbb{C}^{n}_{*}:G_{t}(z)=0,\|z\|\leq r\}$. Moreover, it also restricts as diffeomorphisms
    $\psi_{t}:(\mathbb{S}_{r}^{2n-1},\partial E_{t}(r))\longrightarrow(\mathbb{S}_{r}^{2n-1},\partial E_{1}(r)),$
@@
In the first item, for $t=0$ and $t=1$ the assertion is already proved. In the other cases, it is enough to notice that the Newton polyhedron of $g_{t}^{i}$ is the polyhedron of a complex Pham-Brieskorn polynomial and we reduce to a case already settled. For the second item, by Remark 2.21, the variety $V_{t}^{i}$ intersects $\mathbb{S}^{2n-1}_{r}$ transversely, and thus the same assertion holds for $V_{t}=V_{t}^{1}\cap\dots\cap V_{t}^{k}$. The last statement follows from Ehresmann fibration theorem for subbundles applied to the canonical projections
@@
Remark 2.23.
Let $V\subset(\mathbb{C}^{n},0)$ be a real analytic variety with an isolated singularity at the origin. Then $V\pitchfork\mathbb{S}_{r}^{2n-1}$ for every sufficiently small $r>0$. However, in the previous proof one cannot ensure the same range of $r$ for all $t\in[0,1]$. This condition is called Milnor radius stability in [27]. Therefore, the assertion in Remark 2.21 is essential.
@@
Let $F$ be a complex Hamm map and $G_{t}$ its deformation as above. For each $r>0$ and $0\leq t\leq 1$ fixed, the following statements hold.
@@
    $\psi_{t}:(\mathbb{S}_{r}^{2n-1},K_{G_{t}})\longrightarrow(\mathbb{S}_{r}^{2n-1},K_{F}),$
    where $K_{G_{t}}$ and $K_{F}$ are the links defined by $F$ and $G_{t}$, respectively.
@@
    Let $I\subset\{1,\dots,n\}$ such that $|I|>k$. Then the map $\psi_{t}$ also restricts to a diffeomorphism
    $\psi^{I}_{t}:(\mathbb{S}_{r}^{2|I|-1},K_{G^{I}_{t}})\longrightarrow(\mathbb{S}_{r}^{2|I|-1},K_{F^{I}}),$
    where $K_{G^{I}_{t}}$ and $K_{F^{I}}$ are the links defined by the restrictions $F^{I}$ and $G^{I}_{t}$, respectively.
Example 2.25.
Based on the construction of mixed Hamm-ICIS in the previous subsection, we present an example of a non-degenerate and strongly convenient mixed map germ $F=(f^{1},\dots,f^{k}):(\mathbb{C}^{n}\times\mathbb{C}^{k},0)\longrightarrow(\mathbb{C}^{k},0)$. We observe that the associated Newton polyhedron has a part similar to that of a mixed Hamm map, but it is more general. Let $\textbf{a}=(a_{1},\dots,a_{n})$ and $\textbf{b}=(b_{1},\dots,b_{n})$ be vectors of positive integers, such that $a_{i}\neq b_{i}$ for all $i$. As before, we choose a $n\times k$ matrix $\Lambda=(\lambda_{ij})$ in which every $k\times k$-minor is nonzero and define:
$f^{j}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz},w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=f^{j}_{\textbf{a},\textbf{b}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})+w_{j}^{c_{j}+d_{j}}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{d_{j}}+T^{j}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz},w,{}\mkern 3.0mu\overline{\mkern-3.0muw}),$
where $f^{j}_{\textbf{a},\textbf{b}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})=\sum_{i=1}^{n}\lambda_{ij}z_{i}^{a_{i}+b_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}^{b_{i}}$, $c_{j},d_{j}\geq 1$ are integers, and
$T^{j}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz},w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=\sum_{i=1}^{n}\delta_{j}^{i}z_{j}^{\alpha_{ij}+\beta_{ij}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{j}^{\beta_{ij}}w_{j}^{x_{ij}+y_{ij}}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{y_{ij}}h_{j}^{i}(w,{}\mkern 3.0mu\overline{\mkern-3.0muw}),$
with $\delta_{j}^{i}=0,1$, $\delta_{j}^{i}\neq 0$ for some $i$, $\alpha_{ij}>a_{j},\beta_{ij}>b_{j}$, and $x_{ij},y_{ij}>0$ for all $i,j$, and $h_{j}^{i}:(\mathbb{C}^{k},0)\longrightarrow(\mathbb{C},0)$ is a mixed function. We further require one condition more: all $l\times l$ minors of $\Lambda$ are nonzero for every $l\leq k$.
We first prove the convenience. Note that $f^{j}$ is not a convenient mixed function, due to the term $T^{j}$. Let $I\subset\{1,\dots,n+k\}$ such that $|I|>k$. This implies that we set zero to at most n1n-1 coordinates. Thus, $z_{i}\neq 0$ for some $i$ and each $f^{j}$ has a monomial of the form $\lambda_{ij}z_{i}^{a_{i}+b_{i}}{}\mkern 3.0mu\overline{\mkern-3.0muz}_{i}^{b_{i}}$, then $f^{j,I}\not\equiv 0$ for every $j$. Let $|I|=k$, that is, we set zero to at least $n$ coordinates. If $z=0$, then $V_{F}\cap\mathbb{C}^{I}=\{0\}$. Thus, it remains to verify the case in which $w_{i}=0$ for some $i$. We denote by $\mathbb{C}^{n}_{z}$ and $\mathbb{C}^{k}_{w}$ the subspaces of $\mathbb{C}^{n+k}$ associated with the coordinates $z$ and $w$, respectively.
In this case, $l:=|I^{C}\cap\mathbb{C}^{k}_{w}|$ satisfies $l>0$. Let us form the map $F^{I}_{z}:(\mathbb{C}^{m}_{z},0)\longrightarrow(\mathbb{C}^{l},0)$ whose coordinate functions are those $f^{j}$ for which $w_{j}=0$, and then this map depends only on $(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})$. Notice that:
$|I\cap\mathbb{C}^{k}_{w}|+|I^{C}\cap\mathbb{C}^{k}_{w}|$ $=k,$
$|I\cap\mathbb{C}^{n}_{z}|+|I^{C}\cap\mathbb{C}^{n}_{z}|$ $=n.$
This leads to $m=l$. Since $F^{I}_{z}$ is a mixed Hamm map, the condition on $l\times l$ minors implies that it is non-degenerate, and $F^{I}_{z}=0$ if and only if $z=0$. As we have seen, in this case, $w=0$ and we prove that $F$ is convenient.
The choice of exponents in the mixed functions $T^{j}$ gives that the Newton polyhedron of the coordinate functions of $F$ is the same as that of the map $G:(\mathbb{C}^{n}\times\mathbb{C}^{k},0)\longrightarrow(\mathbb{C}^{k},0)$ whose coordinate functions are:
$g^{j}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz},w,{}\mkern 3.0mu\overline{\mkern-3.0muw})=f^{j}_{\textbf{a},\textbf{b}}(z,{}\mkern 3.0mu\overline{\mkern-3.0muz})+w_{j}^{c_{j}+d_{j}}{}\mkern 3.0mu\overline{\mkern-3.0muw}_{j}^{c_{j}}.$
One can see $G$ as the pullback of a linear holomorphic map by a mixed covering. The linear map $H:(\mathbb{C}^{n}\times\mathbb{C}^{k},0)\longrightarrow(\mathbb{C}^{k},0)$ has coordinate functions $H^{j}=\sum_{i=1}^{k}\lambda_{ij}z_{i}+w_{j}$. The face maps are restrictions of $G$ to subspaces of $\mathbb{C}^{n+k}$. If $|I|=k$, then the same argument shows that $V_{H^{I}}\cap\mathbb{C}^{I}=\{0\}$. If $|I|>k$ and $|I\cap\mathbb{C}_{z}^{n}|>k$, we are in the situation of the mixed Hamm ICIS, which is non-degenerate. If $|I\cap\mathbb{C}_{z}^{n}|=l<k$ and $|I\cap\mathbb{C}_{w}^{k}|=m$ with $l+m>k$, then the diagonal block of the Jacobian matrix of $H^{I}$ related to $w$ has rank $m$. Since the $l\times l$ minors of $\Lambda=(\lambda_{ij})$ are non-vanishing, up to a coordinate change, $J_{H^{I}}$ is decomposed into two blocks that form a matrix of rank $k\times k$. This proves the non-degeneracy of $G$ by the third item of Proposition 2.9.
2.3. Milnor sets